How old is Anna if she's 3yrs older than chuck, and in 5yrs, her age and chucks age would equal 54?
Answer by lwsshak3(11628)

How old is Anna if she's 3yrs older than chuck, and in 5yrs, her age and chucks age would equal 54?
**
let x=anna's present age
x-3=chuck's present age
..
in 5 yrs
x+5=anna's age
x-3+5=x+2=chuck's age
..
(x+5)+(x+2)=54
2x+7=54
2x=47
x=23.5
anna's present age=23.5